Arberg Company’s controller prepared the following budgeted income statement for the coming year:

Sales $423,000
Variable cost 329,940
Contribution margin $93,060
Fixed cost 51,040
Operating income $42,020
Required:
1. What is Arberg’s variable cost ratio? What is its contribution margin ratio?
2. Suppose Arberg’s actual revenues are $30,100 more than budgeted. By how much will operating income increase? Give the answer without preparing a new income statement.
3. How much sales revenue must Arberg earn to break even? Prepare a contribution margin income statement to verify the accuracy of your answer.
4. What is Arberg’s expected margin of safety?
5. What is Arberg’s margin of safety if sales revenue is $384,000?
Answer #1

1.

Variable cost ratio = Variable cost/Sales

= 329,940/423,000

= 78%

Contribution margin ratio = Contribution margin/Sales

= 93,060/423,000

= 22%

2.

Operating income = Sales x Contribution margin ratio – Fixed costs

= 453,100 x 22% - 51,040

= $48,642

Hence, operating income will increase by = 48,642 - 42,020

= $6,622

3.

Break even point ($) = Fixed cost/Contribution margin ratio

= 51,040/22%

= $232,000

Contribution income statement

Sales 232,000
Variable cost - 180,960
Contribution margin 51,040
Fixed cost - 51,040
Operating income 0

4.

Margin of safety = Present sales - Break even sales

= 423,000 - 232,000

= $191,000

5.

Margin of safety = Present sales - Break even sales

= 384,000 - 232,000

= $152,000